crwで JA Avy bal.nu , what iS Tut ett uta mta
Answer #1
A B C D E F G H I J K L M
2
3 A compensating balance is the balance that must be kept with bank in order to take loan from the bank.
4 Thus the available balance of the loan is calculated by deducting the compensatory balance from the total loan amount.
5
6 Total Loan Amount $475,000
7 Period 180 days
8 APR 7.50%
9
10 Nominal interest =Loan Amount*APR*Period (in days)/360
11 $17,812.50 =D6*D8*D7/360
12
13 Compensating Balance % 15%
14 Compensating Balance $71,250 =D6*D13
15
16 Available Balance =Total loan amount - Compensatory Balance
17 $403,750 =D6-D14
18
19 Thus the interest is $17,812.50 on net cash borrowed of $403,750 for 180 days.
20
21 Effective rate =(Nominal interest / Available Balance)*(360 / Period in Days)
22 8.82% =(D11/D17)*(360/D7)
23
24 Hence Effective rate is 8.82%
25